OpenStudy (pulsified333):

For what numbers a and b is the system of equations below satisfied by x=b and y=a? x-5y=a 2x+2y=42 b = a =

OpenStudy (anonymous):

if i read this correctly that means you are to solve \[b-5a=a\\ 2a+2b=42\]

OpenStudy (pulsified333):

thats where I messed up hahaha

OpenStudy (anonymous):

the first can be written as \(b=6a\) and the second as \(a+b=21\)

OpenStudy (anonymous):

solve \(a+6a=21\) in your head, then find \(b\)

OpenStudy (pulsified333):

ok

OpenStudy (pulsified333):

a=3 b=18?